OpenStudy (anonymous):

Suppose you are offered just one slice from a round pizza (in other words, a sector of a circle)The slice must have a perimeter of 32 inches. What diameter pizza will reward you with the largest slice? Specify the diameter and the maximum area of the slice.

imqwerty (imqwerty):

 remember that \[l=r \theta\] and area of sector/slice is given by-\[Area=r^2 \frac{ \theta }{ 2 }\] perimeter=2r+l 32=2r+rθ (32-2r)/r =θ put this theta in that area equation\[area=r^2 \frac{ \theta }{ 2 }\] \[area=r^2 \frac{ (32-2r) }{ r }\]\[area=r(32-2r)\]\[area=32r-2r^2\]now we differentiate this equation with respect to r doing this we will get the value of x for which area will be maximum \[0=32-4r\]\[r=8\] so when radius=8 we get the max area we are asked the diameter diameter=2x radius diameter=2x8=16 we are also asked for area of slice we got this-\[area=32r-2r^2\] so just plug r=8 we get this- area=128